OpenStudy (anonymous):

If the APR of a savings account is 3.6% and interest is compounded monthly, what is the approximate APY of the account?

OpenStudy (anonymous):

someone help me pls

ganeshie8 (ganeshie8):

APY = \(\large \left(1 + \frac{r}{n}\right)^{n} - 1\)

ganeshie8 (ganeshie8):

r = 0.036 n = 12

ganeshie8 (ganeshie8):

plug and chug

ganeshie8 (ganeshie8):

APY = \(\large \left(1 + \frac{0.036}{12}\right)^{12} - 1\)

ganeshie8 (ganeshie8):

use ur calculator

OpenStudy (anonymous):

my calculator is a phone so i cant use it is the answer 3.36

ganeshie8 (ganeshie8):

APY = \(\color{red}{3.66}\)

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

thank you

ganeshie8 (ganeshie8):

u wlc ^_^

OpenStudy (anonymous):

Apr is 3.4% intrest compounded semianually no additional deposits or withdraws for an entire year what will the balance of the acount be after all the intrest is paid for a year on a principal balance of $9700

OpenStudy (anonymous):

help pls

OpenStudy (anonymous):

someone?

ganeshie8 (ganeshie8):

use compound interest formula : \(A = P\left(1 + \dfrac{r}{n}\right)^{nt}\)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

still can't get it

ganeshie8 (ganeshie8):

figure out the numbers and plugin

OpenStudy (anonymous):

wait is it 9733

OpenStudy (anonymous):

or is it 9732

OpenStudy (anonymous):

is it

ganeshie8 (ganeshie8):

no dude P = 9700 r = 0.034 n = semiannually = 2

ganeshie8 (ganeshie8):

\(A = 9700\left(1 + \dfrac{0.034}{2}\right)^{2*1}\)

ganeshie8 (ganeshie8):

use wolfram : http://www.wolframalpha.com/input/?i=9700%281+%2B+0.034%2F2%29%5E%282%29